# A deep backward regression-based scheme for high-dimensional nonlinear partial differential equations

### 16 Mar 2026 v1

@@ 5 Numerical experiments
However, a markedly different pattern emerges as the dimension increases beyond $10$. For $d=15$ (Table 5), the DBDP1 method exhibits a dramatic deterioration in performance, with relative errors soaring to approximately $61.5357\%$, $44.5369\%$ and $55.5073\%$ for $N=120,130,140$, respectively. These unacceptably large errors indicate that DBDP1 essentially fails to provide reliable approximations in this higher-dimensional setting, a finding that corroborates the observations in [28] where the authors noted that DBDP1 becomes ineffective when the dimension exceeds 10. In stark contrast, the DBR method maintains relative errors consistently below $10\%$ across all tested temporal discretizations, achieving an impressive $1.2529\%$ relative error at $N=140$. This remarkable stability and accuracy demonstrate the robustness and superior approximation capability of DBR in high-dimensional regimes.
The advantage of DBR becomes even more pronounced at $d=20$ (Table 6). While DBDP1 produces relative errors of $70.3907\%$, $54.9322\%$ and $80.2554\%$ for $N=90,100,140$, rendering its estimates practically meaningless, DBR achieves remarkably low relative errors of $6974\%, 3.3$$3720\%$ and $9.0293\%$ under the same conditions. This striking performance gap unequivocally demonstrates that the DBR algorithm retains its accuracy and stability in high-dimensional regimes where traditional methods like DBDP1 completely break down.
These numerical results collectively substantiate the core contribution of this work: by replacing the pathwise loss function with expectation-based alternatives, the DBR method effectively mitigates the adverse effects of increasing dimensionality, thereby offering a reliable and accurate approach for solving high-dimensional nonlinear PDEs. The method’s consistent performance across dimensions $d=15$ and $d=20$, particularly when compared to the catastrophic failure of DBDP1, provides compelling evidence of its practical utility for real-world high-dimensional problems. Due to prohibitive computational costs, the high-dimensional regime ($d>20$) was excluded from our experiments. We anticipate that the results would benefit from an increased sample size.
@@ 6 Conclusion
In this paper, we have proposed a new deep probabilistic numerical algorithm which is called deep backward regression-based method for high-dimensional quasi-linear PDEs based on the probabilistic representation of the solution of the PDEs. The computation of our loss functions essentially involves performing a local Monte Carlo simulation at each time step, using the average value to replace the individual sample value for calculating the labels. Its advantage is that we do not need to fit the Brownian motion (Since the noise is not fittable, this wastes the capacity of the algorithm.) and we merely need to fit the average dynamics of the system. Thus, our DBR method has a higher learning efficiency. Then we have rigorously proven that the error estimates of the proposed DBR method. Moreover, we have established the half-order convergence of our deep numerical scheme by rigorous analysis. Furthermore, it is interesting and challenging that one develops an optimization of sampling strategy to reduce the computational cost of conditional expectations or an lightweight the neural network architecture to reduce computational cost of the models to improve the the numeral method of PDEs.
@@
Data sharing is not applicable to this article as no datasets are generated or analyzed in this study.

### 21 May 2026 v3

@@ 5 Numerical experiments
However, a markedly different pattern emerges on this benchmark when the dimension is increased to $d=15$. As shown in Table 5.5, the DBDP1 method exhibits a substantial deterioration in performance, with relative errors of approximately $61.5357\%$, $44.5369\%$, and $55.5073\%$ for $N=120,130,140$, respectively. These large errors indicate that DBDP1 is less reliable for this specific challenging unbounded example at this dimension, rather than suggesting a universal failure of DBDP1 for all problems with dimension larger than ten. In contrast, the DBR method maintains relative errors below $10\%$ across the tested temporal discretizations, and achieves a relative error of $1.2529\%$ at $N=140$. This comparison highlights the robustness of the DBR scheme for the present benchmark.
The performance gap becomes even more pronounced at $d=20$ (Table 5.6). DBDP1 produces relative errors of $70.3907\%$, $54.9322\%$, and $80.2554\%$ for $N=90,100,140$, respectively, whereas DBR achieves relative errors of $6974\%, 3.3$$3720\%$, and $9.0293\%$ under the same conditions. These results show that, for this particular unbounded benchmark, the expectation-based DBR loss leads to more stable and accurate estimates than the pathwise DBDP1 loss in the tested high-dimensional settings.
These numerical results collectively support the core contribution of this work: by replacing the pathwise loss function with expectation-based alternatives, the DBR method can effectively reduce the adverse impact of pathwise noise and improve training stability. For the challenging unbounded benchmark considered in Example 2, the method’s performance at $d=15$ and $d=20$, particularly when compared with the substantial deterioration of DBDP1 on the same benchmark, provides evidence of its practical utility for a class of high-dimensional nonlinear PDEs with complex and unbounded structures. Due to prohibitive computational costs, the high-dimensional regime $d>20$ was excluded from our experiments. We anticipate that the results would benefit from an increased sample size.
@@ 6 Conclusion
In this paper, we proposed a deep backward regression-based method for solving high-dimensional quasilinear parabolic PDEs through their FBSDE representations. The main idea is to replace pathwise Euler residuals by conditional-expectation-based training targets. These targets are approximated through local multi-path Monte Carlo averaging, which reduces the influence of Brownian pathwise noise and yields a smoother regression problem for the neural networks.
Under the population-loss convention and exact minimization setting, we established error estimates for the proposed DBR method. Under suitable approximation and integrability assumptions, we further proved a half-order convergence result. We also extended the method to variational inequalities and derived a corresponding error estimate for the reflected case.
Several directions remain open. First, more efficient sampling strategies for approximating the conditional expectations could reduce the computational cost of the method. Second, lighter neural-network architectures may improve scalability in very high dimensions. Finally, it would be valuable to incorporate finite-sample generalization errors and optimization errors into the theoretical analysis, thereby bridging the gap between the idealized population-loss theory and practical SGD training.
@@
All data used in the numerical experiments (Examples 1-2) are generated procedurally within the code using the parameters defined in the paper.